# QZO: A Catalog of 5 Million Quasars from the Zwicky Transient Facility

### v1

@@ Abstract
Machine learning methods are well established in the classification of quasars (QSOs). However, the advent of light curve observations adds a great amount of complexity to the problem. Our goal is to use the Zwicky Transient Facility (ZTF) to create a catalog of QSOs. We process the ZTF DR20 light curves with a transformer artificial neural network and combine the Pan-STARRS (PS), AllWISE, and Gaia surveys with extreme gradient boosting. Using ZTF g-band data with at least 100 observational epochs per light curve, we obtain 97% F1 score for QSOs. We find that with 3 day median cadence, a survey time span of at least 900 days is required to achieve 90% QSO F1 score. However, one can obtain the same score with a survey time span of 1800 days and the median cadence prolonged to 12 days. We find that ZTF classification is superior to the PS static bands, and on par with WISE and Gaia measurements. Additionally, we find that the light curves provide the most important features for QSO classification in the ZTF dataset. We robustly classify objects fainter than the $5\sigma$ SNR limit at $g=20.8$ by requiring $g<\mathrm{n_{obs}}/80+20.375$. For this sample, we run inference with added WISE observations, and find 4,849,574 objects classified as QSOs. For 33% of QZO objects, with available WISE data, we publish redshifts with estimated error $\Delta z/(1+z)=0.14$. We release the catalog at: placeholder [^1] Data will be published upon acceptance in a journal., and the code at: https://github.com/snakoneczny/ztf-agn.
@@
facilities: Zwicky Transient Facility (ZTF, Bellm et al., 2019a; Graham et al., 2019; Masci et al., 2019; Dekany et al., 2020), Sloan Digital Sky Survey (SDSS, Almeida et al., 2023), Pan-STARRS (PS, Chambers et al., 2016), Wide-Field Infrared Survey Explorer (WISE, Wright et al., 2010; Mainzer et al., 2011; Cutri et al., 2021), Gaia (Gaia Collaboration et al., 2016, 2021)software: Python 3 (Van Rossum & Drake, 2009), Tensorflow (Abadi et al., 2015), Keras (Chollet et al., 2015), Astromer (Donoso-Oliva et al., 2023), XGBoost (Chen & Guestrin, 2016), Scikit-learn (Pedregosa et al., 2011), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), SkyPortal (van der Walt et al., 2019; Coughlin et al., 2023), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), IPython (Pérez & Granger, 2007), Pandas (Wes McKinney, 2010), Matplotlib (Hunter, 2007), seaborn (Waskom, 2021), tqdm (da Costa-Luis, 2019)
@@ 1 Introduction
Quasars are a type of active galaxy, usually occupying massive dark matter halos (Eftekharzadeh et al., 2015; DiPompeo et al., 2016), and emitting enormous amounts of energy through accretion of matter onto a supermassive black hole (Kormendy & Ho, 2013). They mostly appear as faint and point like objects. Due to their highly biased tracing of vast volumes of the large scale structure (DiPompeo et al., 2014; Laurent et al., 2017), quasars are used for various cosmological applications: a test of the cosmological principle of isotropy and homogeneity (Secrest et al., 2021; Dam et al., 2023), the growth rate of structure (García-García et al., 2021; Alonso et al., 2023), primordial non-Gaussianity (Leistedt et al., 2014; Castorina et al., 2019; Krolewski et al., 2024), the Hubble distance (Hou et al., 2020), baryon acoustic oscillations (Ata et al., 2017; Zarrouk et al., 2021), the integrated Sachs-Wolfe effect (Stölzner et al., 2018), the expansion rate of the universe as standardizable candles (Setti & Woltjer, 1973; Risaliti & Lusso, 2015; Lusso et al., 2020), calibration of the reference frames for Galactic studies (Lindegren et al., 2018), measurement of halo masses (DiPompeo et al., 2017), cross-correlation with the $\gamma$-ray background (Cuoco et al., 2017), tomographic angular clustering (Leistedt et al., 2014; Ho et al., 2015), quasar bias measurements (Sherwin et al., 2012), the baryon density (Yahata et al., 2005), cosmic magnification (Scranton et al., 2005), and cross correlation with foreground galaxies as a probe of weak lensing (Ménard & Bartelmann, 2002).
@@
Quasars are also variable objects well described by a damped random walk (Kelly et al., 2009; Kozłowski et al., 2010; MacLeod et al., 2010). However, there are also fluctuations in a form of, e.g. tidal disruption events (TDEs, Rees, 1988; Gezari, 2021; Stein et al., 2024), and possibly gravitational wave (GW) flares (McKernan et al., 2019; Graham et al., 2020; Wang et al., 2021; Kimura et al., 2021; Perna et al., 2021). The latter have not yet been confirmed, but Graham et al. (2023) present nine candidate counterparts for such events from Ligo/Virgo detectors (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015). The importance of such events results from accretion disks being the only channel for stellar mass black hole mergers which reliably produce electromagnetic (EM) counterparts, and thus the identification of GW event with a host galaxy at a known redshift (Ashton et al., 2021; Calderón Bustillo et al., 2021; Palmese et al., 2021). The GW signal then becomes a standard siren, which can be used to measure the expansion history of the universe (Mukherjee et al., 2020; Chen et al., 2022). However, detection of EM counterparts in AGN (active galactic nucleus) time series is challenging due to potentially too bright or too thick accretion disks, as well as the incomplete census of AGN. Hence, a first step in a systematic search of such flares is a construction of AGN catalog with available time-series data.
@@
The most reliable source of quasar [^2] We use AGN and QSO terms interchangeably. identification and redshifts are galaxy spectra. Spectroscopic surveys have provided $\sim 10^{5}$-$10^{6}$ QSOs, e.g. 2dF QSO Redshift Survey (2QZ, Croom et al., 2004), 2dF-SDSS LRG and QSO (2SLAQ, Croom et al., 2009), and the Sloan Digital Sky Survey (SDSS, York et al., 2000; Lyke et al., 2020). Ongoing and future spectroscopic surveys such as DESI (DESI Collaboration et al., 2016) and 4MOST (de Jong et al., 2019; Merloni et al., 2019; Richard et al., 2019) expect to obtain spectra for 3 million quasars. Since QSOs are sparsely distributed in comparison to inactive galaxies, wide area photometric surveys are a much better way of observing large numbers of QSOs. However, the lack of spectra makes classification more complicated and redshifts much less accurate. Nearly 3 million quasars with photometric redshifts (photo-zs) have been cataloged (e.g. Shu et al., 2019; Kunsági-Máté et al., 2022; Storey-Fisher et al., 2024), through surveys such as the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010), unWISE (Lang, 2014), Pan-STARRS (Chambers et al., 2016), and Gaia (Gaia Collaboration et al., 2016). In the future, the Rubin Observatory’s LSST will photometrically observe up to 10 million quasars (Ivezić, 2016).
@@
One possibility to classify objects in photometric data is SED fitting, which depending on available data can derive physical properties (Ciesla et al., 2015; Stalevski et al., 2016; Calistro Rivera et al., 2016; Yang et al., 2020; Małek et al., 2020), as well as photo-zs (Salvato et al., 2009, 2011; Fotopoulou et al., 2016; Fotopoulou & Paltani, 2018). Also, using different magnitudes, one can create a color-color space which allows for object classification (Warren et al., 2000; Maddox et al., 2008; Edelson & Malkan, 2012; Stern et al., 2012; Wu et al., 2012; Secrest et al., 2015; Assef et al., 2018). More sophisticated methods include probabilistic models (Richards et al., 2004, 2009a, 2009b; Bovy et al., 2011, 2012; DiPompeo et al., 2015; Richards et al., 2015), but due to the high complexity of multidimensional color-color space, machine learning (ML) methods achieve the best results (Brescia et al., 2015; Carrasco et al., 2015; Kurcz et al., 2016; Nakoneczny et al., 2019; Logan & Fotopoulou, 2020). Depending on the availability of infrared observations, one can also estimate photo-zs with ML (Brescia et al., 2013; Yang et al., 2017; Pasquet-Itam & Pasquet, 2018; Curran, 2020; Nakoneczny et al., 2021).
@@
Another type of photometric observation are time series data, provided by surveys such as ASAS (Pojmanski, 2002), NSVS (Woźniak et al., 2004; Hoffman et al., 2009), the Palomar Transient Factory (Law et al., 2009), the Catalina Surveys (Drake et al., 2014, 2017), and ASAS-SN (Kochanek et al., 2017; Jayasinghe et al., 2018). The Zwicky Transient Facility (ZTF, Bellm et al., 2019a; Graham et al., 2019; Masci et al., 2019; Dekany et al., 2020) has operated since March 2018 and covers the northern sky with 3 day cadence. Within the time-series data, ML was applied to classify microlensing events (Godines et al., 2019), transients (Stachie et al., 2020; Gomez et al., 2023; Rehemtulla & Miller, 2023), and variables (Richards et al., 2011; García-Jara et al., 2022; Mistry et al., 2022). The ZTF Source Classification Project (SCoPe, van Roestel et al., 2021; Coughlin et al., 2021; Healy et al., 2024) provides classification for many different classes of variables. However, our goal is to create a method fully optimized for quasar detection, apply deep learning methods, and achieve the highest possible accuracy. Observations of quasar light curves are relatively new, and their usage in classification has not yet been fully established in comparison to other astronomical observations. ZTF is well suited for this goal, as it delivers light curves with irregular sampling, over a six year period of observations.
@@
Our first goal is to test quasar classification with deep learning methods based on the transformer architecture, and compare the results with shallow feature extraction methods. Next, we establish the minimum requirements on the length and sampling of time-series data in order to achieve reliable quasar classification. We compare ZTF classification to visible griz bands from Pan-STARRS DR1 (PS, Chambers et al., 2016), mid-infrared W[1-4] filters from AllWISE (Wright et al., 2010; Mainzer et al., 2011; Cutri et al., 2021), and parallax and proper motion measurements from Gaia EDR3 (Gaia Collaboration et al., 2016, 2021). Finally, we use these results to create a high quality quasar catalog based on the ZTF data.
@@ 2.2 Inference data
We perform the inference using ZTF data release (DR) 20, covering March 2018 to January 2024. For inference, we use only the g-band data, while we add the r-band to experiments, and omit the i-band due to its much lower coverage compared to the two other bands. In total, ZTF DR20 consists of 1.5B and 2.4B objects in the g- and r-bands, respectively. The estimated number of bad epochs per light curve, with likely suspect and unusable photometry, is 8% and 11% in g and r, respectively [^3] See Table 1 in https://irsa.ipac.caltech.edu/data/ZTF/docs/releases/dr20/ztf_release_notes_dr20.pdf. We remove deep drilling observations by averaging observations within the same night. Finally, after removal of light curves with less than 20 observations, we obtain the inference dataset of 789M objects in the ZTF g band. Additionally, we either mark or remove duplicated light curves by finding objects which within $1\arcsec$ have a neighbor with more observation epochs. The number of deduplicated light curves equals 533M.
@@ 2.3 Training data
We obtain the training data in g and r bands by cross-matching the ZTF data with SDSS DR18 (SDSS, Almeida et al., 2023), using a $1\arcsec$ distance threshold, and applying the same preprocessing as in the case of inference data. We clean the SDSS data by removing objects with known redshift estimation errors as encoded in the zWarning flag, with the exception of the fifth bit of this flag, which does not signal problems. Additionally, we remove duplicated ZTF light curves by accepting only the longest light curve among the neighbors within the $1\arcsec$ distance. We analyze both g- and r-band datasets independently. The training set size is 483k and 665k in g- and r-bands. The class distribution is 47% galaxies, 25% quasars, and 28% stars in the g-band, and 59% galaxies, 19% quasars, and 22% stars in the r-band. It totals to 645062 and 703685 quasars in g- and r-bands, respectively. We compute the experiments with shallow precomputed features (Healy et al., 2024) in Section 4.1, based on ZTF DR5, published in January 2021, since at the time of writing, shallow features were only computed for this DR. However, we did not first utilize all the light curves from this dataset, and could not update it later due to removal of data from the servers. Hence, the DR5 training dataset size is 301k and 381k in g- and r-bands, respectively, which also lowers the deep learning results due to a smaller training set size. We use the DR5 data only for comparison between the shallow and deep learning results of light curve classification. We use 37 of the precomputed features, omitting the magnitude-time histograms and ZTF alerts information (see Table 1 in Healy et al., 2024).
Figure 1 shows distributions of median magnitude and number of observations for inference and training data, also separated for different classes. The plots are normalized to unity, and the class plots are divided by three in order to properly compare between the different classes. We see on the magnitude plot that the training data cover the inference data, but the training distribution is different from the inference one, which together with the class distribution, hint to a larger fraction of quasars present at the faint end of inference data. Due to this, the final classification of inference data requires additional testing, and the results will not be calibrated, meaning that among objects with estimated QSO probability around 90%, the actual number of quasars will be different than the 90%. The number of observations plot shows that there are no significant differences between the classes. Such differences could constitute a data leak, as it could be used as additional information for classification, which in fact might not be the same in the inference data.

### v2

@@ Abstract
Machine learning methods are well established in the classification of quasars (QSOs). However, the advent of light curve observations adds a great amount of complexity to the problem. Our goal is to use the Zwicky Transient Facility (ZTF) to create a catalog of QSOs. We process the ZTF DR20 light curves with a transformer artificial neural network and combine different surveys with extreme gradient boosting. Based on ZTF g-band and WISE observations, we find 4,849,574 objects classified as QSOs with confidence higher than 90%. We robustly classify objects fainter than the $5\sigma$ SNR limit at $g=20.8$ by requiring $g<n_{\mathrm{obs}}/80+20.375$. For 33% of QZO objects, with available WISE data, we publish redshifts with estimated error $\Delta z/(1+z)=0.14$. We find that ZTF classification is superior to the Pan-STARRS static bands, and on par with WISE and Gaia measurements, but the light curves provide the most important features for QSO classification in the ZTF dataset. Using ZTF g-band data with at least 100 observational epochs per light curve, we obtain 97% F1 score for QSOs. We find that with 3 day median cadence, a survey time span of at least 900 days is required to achieve 90% QSO F1 score. However, one can obtain the same score with a survey time span of 1800 days and the median cadence prolonged to 12 days. We release the catalog, models and code on Zenodo [^1] https://zenodo.org/records/16410988222https://zenodo.org/records/16535608, and GitHub [^3] https://github.com/snakoneczny/ztf-agn.
@@
facilities: Zwicky Transient Facility (ZTF, Bellm et al., 2019a; Graham et al., 2019; Masci et al., 2019; Dekany et al., 2020), Sloan Digital Sky Survey (SDSS, Almeida et al., 2023), Pan-STARRS (PS, Chambers et al., 2016), Wide-Field Infrared Survey Explorer (WISE, Wright et al., 2010; Cutri et al., 2021), Gaia (Gaia Collaboration et al., 2016, 2021)software: Python 3 (Van Rossum & Drake, 2009), Tensorflow (Abadi et al., 2015), Keras (Chollet et al., 2015), Astromer (Donoso-Oliva et al., 2023), XGBoost (Chen & Guestrin, 2016), Scikit-learn (Pedregosa et al., 2011), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), SkyPortal (van der Walt et al., 2019; Coughlin et al., 2023), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), IPython (Pérez & Granger, 2007), Pandas (Wes McKinney, 2010), Matplotlib (Hunter, 2007), seaborn (Waskom, 2021), tqdm (da Costa-Luis, 2019)
@@ 1 Introduction
Quasars are a type of active galaxy, usually occupying massive dark matter halos (Eftekharzadeh et al., 2015; DiPompeo et al., 2016), and emitting enormous amounts of energy through accretion of matter onto a supermassive black hole (Kormendy & Ho, 2013). They mostly appear as faint and point like objects. Due to their highly biased tracing of vast volumes of the large scale structure (DiPompeo et al., 2014; Laurent et al., 2017), quasars are used for various cosmological applications: a test of the cosmological principle of isotropy and homogeneity (Secrest et al., 2021; Dam et al., 2023), the growth rate of structure (Garc´ıa-Garc´ıa et al., 2021; Alonso et al., 2023), primordial non-Gaussianity (Leistedt et al., 2014; Castorina et al., 2019; Krolewski et al., 2024), the Hubble distance (Hou et al., 2020), baryon acoustic oscillations (Ata et al., 2017; Zarrouk et al., 2021), the integrated Sachs-Wolfe effect (Stölzner et al., 2018), the expansion rate of the universe as standardizable candles (Setti & Woltjer, 1973; Risaliti & Lusso, 2015; Lusso et al., 2020), calibration of the reference frames for Galactic studies (Lindegren et al., 2018), measurement of halo masses (DiPompeo et al., 2017), cross-correlation with the $\gamma$-ray background (Cuoco et al., 2017), tomographic angular clustering (Leistedt et al., 2014; Ho et al., 2015), quasar bias measurements (Sherwin et al., 2012), the baryon density (Yahata et al., 2005), cosmic magnification (Scranton et al., 2005), and cross correlation with foreground galaxies as a probe of weak lensing (Ménard & Bartelmann, 2002).
Quasars are also variable objects, which can be described by a damped random walk (Kelly et al., 2009; Kozłowski et al., 2010). The model is consistent with the observed light curves on timescales from several months to a few years (MacLeod et al., 2010; Zu et al., 2011; MacLeod et al., 2012), but diverges at scales less than a few months (Mushotzky et al., 2011; Zu et al., 2013; Kasliwal et al., 2015; Stone et al., 2022). The physical source of quasar variability is not well understood (Ulrich et al., 1997; Padovani et al., 2017). Studies suggested that variability amplitude increases with decreasing luminosity, rest frame wavelength, and Eddington ratio (Wills et al., 1993; Giveon et al., 1999; Vanden Berk et al., 2004; Wold et al., 2007), and the characteristic variability time-scale depends on the black hole mass (Collier & Peterson, 2001; Kelly et al., 2009; MacLeod et al., 2010; Simm et al., 2016). Burke et al. (2021) found a strong correlation between the damping time-scale and black hole mass, and (Stone et al., 2022) found a weak wavelength dependence of damping time scale.
@@
There are also fluctuations in a form of, e.g. tidal disruption events (TDEs, Rees, 1988; Gezari, 2021; Stein et al., 2024), and possibly gravitational wave (GW) flares (McKernan et al., 2019; Graham et al., 2020; Wang et al., 2021; Kimura et al., 2021; Perna et al., 2021). The latter have not yet been confirmed, but Graham et al. (2023) present nine candidate counterparts for such events from Ligo/Virgo detectors (LIGO Scientific Collaboration et al., 2015; Acernese et al., 2015). The importance of such events results from accretion disks being the only channel for stellar mass black hole mergers which reliably produce electromagnetic (EM) counterparts, and thus the identification of GW event with a host galaxy at a known redshift (Ashton et al., 2021; Calderón Bustillo et al., 2021; Palmese et al., 2021). The GW signal then becomes a standard siren, which can be used to measure the expansion history of the universe (Mukherjee et al., 2020; Chen et al., 2022). However, detection of EM counterparts in AGN (active galactic nucleus) time series is challenging due to potentially too bright or too thick accretion disks, as well as the incomplete census of AGN. Hence, a first step in a systematic search of such flares is a construction of AGN catalog with available time-series data.
@@
The most reliable source of quasar [^4] We use AGN and QSO terms interchangeably. identification and redshifts are spectra. Spectroscopic surveys have provided $\sim 10^{5}$-$10^{6}$ QSOs, e.g. 2dF QSO Redshift Survey (2QZ, Croom et al., 2004), 2dF-SDSS LRG and QSO (2SLAQ, Croom et al., 2009), and the Sloan Digital Sky Survey (SDSS, York et al., 2000; Lyke et al., 2020). Ongoing and future spectroscopic surveys such as DESI (DESI Collaboration et al., 2016) and 4MOST (de Jong et al., 2019; Merloni et al., 2019; Richard et al., 2019) expect to obtain spectra for 3 million quasars. Since spectroscopic surveys are time-consuming, wide area photometric surveys are a much better way of observing large numbers of QSOs. However, the lack of spectra makes classification more complicated and redshifts much less accurate. Nearly 3 million quasars with photometric redshifts (photo-zs) have been cataloged (e.g. Shu et al., 2019; Kunsági-Máté et al., 2022; Yang & Shen, 2023; Nakazono et al., 2024; Storey-Fisher et al., 2024), through surveys such as the Wide-field Infrared Survey Explorer (WISE, Wright et al., 2010), unWISE (Lang, 2014), Pan-STARRS (Chambers et al., 2016), and Gaia (Gaia Collaboration et al., 2016). In the future, the Rubin Observatory’s LSST will photometrically observe up to 10 million quasars (Ivezić, 2016).
@@
One possibility to classify objects in photometric data is SED fitting, which depending on available data can derive physical properties (Ciesla et al., 2015; Stalevski et al., 2016; Calistro Rivera et al., 2016; Yang et al., 2020; Małek et al., 2020), as well as photo-zs (Salvato et al., 2009, 2011; Fotopoulou et al., 2016; Fotopoulou & Paltani, 2018). Also, using different magnitudes, one can create a color-color space which allows for object classification (Warren et al., 2000; Maddox et al., 2008; Edelson & Malkan, 2012; Stern et al., 2012; Wu et al., 2012; Secrest et al., 2015; Assef et al., 2018). More sophisticated methods include probabilistic models (Richards et al., 2004, 2009a, 2009b; Bovy et al., 2011, 2012; DiPompeo et al., 2015; Richards et al., 2015), but due to the high complexity of multidimensional color-color space, machine learning (ML) methods achieve the best results (Brescia et al., 2015; Carrasco et al., 2015; Kurcz et al., 2016; Nakoneczny et al., 2019; Logan & Fotopoulou, 2020). One can also estimate photo-zs with ML (Brescia et al., 2013; Yang et al., 2017; Pasquet-Itam & Pasquet, 2018; Curran, 2020; Nakoneczny et al., 2021).
@@
Another useful information for quasar classification is the photometric variability across time, provided by surveys such as ASAS (Pojmanski, 2002), NSVS (Woźniak et al., 2004; Hoffman et al., 2009), the Palomar Transient Factory (Law et al., 2009), the Catalina Surveys (Drake et al., 2014, 2017), and ASAS-SN (Kochanek et al., 2017; Jayasinghe et al., 2018). The Zwicky Transient Facility (ZTF, Bellm et al., 2019a; Graham et al., 2019; Masci et al., 2019; Dekany et al., 2020) has operated since March 2018 and covers the northern sky with 3 day cadence. Since the observations of quasar light curves are relatively new, and their usage in classification has not yet been fully established in comparison to other astronomical observations, ZTF is well suited for this goal, as it delivers light curves with irregular sampling, over a six year period of observations.
Several projects in ZTF performed AGN classification alongside other classes of variable and non variable objects. For instance, the ZTF Source Classification Project (SCoPe, van Roestel et al., 2021; Coughlin et al., 2021; Healy et al., 2024) uses convolutional neural networks with time magnitude histograms, and extreme gradient boosting (XGB, Chen & Guestrin, 2016) with precomputed features, trained on a set of $\sim 10^{3}$ manually selected AGNs. Automatic Learning for the Rapid Classification of Events (ALeRCE, Förster et al., 2021; Sánchez-Sáez et al., 2021) classifies the ZTF alert stream, which is a subset of ZTF light curves showing large fluctuation in magnitude, using balanced decision trees with precomputed features, and distinguishing several AGN subclasses. (Sánchez-Sáez et al., 2023) classifies light curves within the overlap of ZTF and 4-m Multi-Object Spectroscopic Telescope (4MOST, de Jong et al., 2019) surveys, distinguishing 3 ranges of redshifts for the AGNs, in order to identify AGN candidates for 4MOST survey. Our goal is to classify all ZTF light curves using a deep learning transformer model trained on SDSS data with 120,000 quasars in our training dataset.
@@
We test quasar classification with deep learning methods based on the transformer architecture, and compare the results with shallow feature extraction methods. Next, we establish the minimum requirements on the length and sampling of time-series data in order to achieve reliable quasar classification. We compare ZTF classification to visible griz bands from Pan-STARRS DR1 (PS, Chambers et al., 2016), mid-infrared W[1-4] filters from AllWISE (Wright et al., 2010; Cutri et al., 2021), and parallax and proper motion measurements from Gaia EDR3 (Gaia Collaboration et al., 2016, 2021). Finally, we use these results to create a high quality quasar catalog based on the ZTF data.
@@ 2.2 Inference data
We perform the inference using ZTF data release (DR) 20, covering March 2018 to January 2024. For inference, we use only the g-band data, while we add the r-band to experiments, and omit the i-band due to its much lower coverage compared to the two other bands. In total, ZTF DR20 consists of 1.5B and 2.4B objects in the g- and r-bands, respectively. The estimated number of bad epochs per light curve, with likely suspect and unusable photometry, is 8% and 11% in g and r, respectively [^5] See Table 1 in https://irsa.ipac.caltech.edu/data/ZTF/docs/releases/dr20/ztf_release_notes_dr20.pdf. We remove deep drilling observations by averaging observations within the same night. Finally, after removal of light curves with less than 20 observations, we obtain the inference dataset of 789M objects in the ZTF g band. Additionally, we either mark or remove duplicated light curves by finding objects which within $1\arcsec$ have a neighbor with more observation epochs. The number of deduplicated light curves equals 533M. In this work, we do not consider morphological features, but these can be used as an extension to our methodology.
@@ 2.3 Training data
We obtain the training data in g and r bands by cross-matching the ZTF data with SDSS DR18 (SDSS, Almeida et al., 2023), using a $1\arcsec$ distance threshold, and applying the same preprocessing as in the case of inference data. We clean the SDSS data by removing objects with known redshift estimation errors as encoded in the zWarning flag, with the exception of the fifth bit of this flag, which does not signal problems. Additionally, we remove duplicated ZTF light curves by accepting only the longest light curve among the neighbors within the $1\arcsec$ distance. We analyze both g- and r-band datasets independently. The training set size is 483k and 665k in g- and r-bands. The class distribution is 47% galaxies, 25% quasars, and 28% stars in the g-band, and 59% galaxies, 19% quasars, and 22% stars in the r-band. It totals to 645062 and 703685 quasars in g- and r-bands, respectively. In Section 4.1, we compute the experiments with shallow precomputed features (see Table 1 in Healy et al., 2024). For this goal, we use ZTF DR5 published in January 2021, since at the time of writing, shallow features were only computed for this DR. However, we did not first utilize all the light curves from this dataset, and could not update it later due to removal of data from the servers. Hence, the DR5 training dataset size is 301k and 381k in g- and r-bands, respectively, which also lowers the deep learning results due to a smaller training set size. We use the DR5 data only for comparison between the shallow and deep learning results of light curve classification. We use 37 of the precomputed features, omitting the magnitude-time histograms and ZTF alerts information.
Figure 1 shows distributions of median magnitude and number of observations for inference and training data, also separated for different classes. The plots are normalized to unity, and the class plots are divided by three in order to properly compare the different classes. We can see a bimodal galaxy magnitude distribution, which does not bias the final inference, as we show in Fig. 17. The training data cover the faint end of the inference data. However, the training distribution is different from the inference one, which together with the class distribution, hint to a larger fraction of quasars present at the faint end of inference data. Due to this, the final classification of inference data requires additional testing, and the results will not be calibrated. It means that if we chose 100 objects from the QSO classification probability range 90% to 91%, the actual number of QSOs in such a sample would be different than 90 or 91 objects, and that would not be due to a chance occurrence. The number of observations plot shows galaxies, quasars, and stars within the whole range of epochs. A shift in these distributions could also bias the final results.